In the recent days we realized that the Green AOM was not configured in double pass. In the morning shift Martina re-configured it in order to have a double pass configuration (see dedicated entry)
At the beginning of the shift the cavity was misaligned and the peaks were very low (1.5V). We managed to align the cavity but the peaks were maximum 4.5 V. Observing the beam shape into the cameras we realized that one of the 2 diaphragram into the green pass was closed. We opened it. The peaks raised at 6 V but the signal for the PDH loop was too low. We realized that the photodiode was misaligned and we realigned it. After that we were able to lock the cavity and we tuned the demodulation phase. In the new configuration the maximum power transimtted by the filter cavity is 6.7V and no more 7.1V.
Observing the two cameras on SQB1 we saw that the SQB1 FI transmitted beam was clipped. Moving SQB1 M21 by about 100 steps in H and in V we recovered it.
We tried to find the TEM00 of the IR in transmission to the cavity but we did not manage to obtain it. We realized that the work on the AOM changed the IR/Green Coresonance into the cavity thus with a scan we found the BAB TEM00 with the AOM offset around 0.8V.
We improved the alignment of the BAB acting on the EQB1 IR Mirrors (M3 and M5). In tranmission from the cavity we have about 0.16V of BAB
Thwn we switched on the subcarrier beam, we found the TEM00 and we improve a little bit its alignment using SC_M1 and SC_M3. The alignment of the SC is not optimal. It is better to run the SC AA (to do it we need to decrease the SC power. We will do it next week).
SQB1 and FCEM SBE recovery:
In the last days the offload caused the saturation of the FCEM_SBE actuators. With the help of Matteo we discharged the filter cavity offload and we recentered the inverted pendulum.
We recentered also the SQB1 inverted pendulum. Now all the corrections of the Suspended benches are low. Thus we can try to have a long lock during the week end.
IR/SC Frequency study:
After the relock of the cavity we found the BAB TEM00 at 0.8030 V of AOM offset, whereas the SC TEM00 was at 0.9305 V of AOM offset. We moved the SC PLL offset in order to have them corensonant into the filter cavity. The initial offset of SC was 100.10057467 MHz, whereas the offset used to obtain the coresonance was 100.09427962 MHz. Thus the AOM scan frequency slope is 49.37 kHz/V. Before the intervention on the AOM was 25 kHz/V. This confirm that now we are in double pass configuration.
As second step we went into the DET e room and we moved the offset of the AOM frequency generator in order to have back the Green/IR coresonance around 0V. The initial frequency was 80.05370 MHz, now the frequency is 80.01433 MHz. From this we confirm again that the AOM frequency slope is 50 kHz/V. Now the ramp range should be enough to scan a big fraction of a free spectral range. In particular we are able to see TEM00, TEM10 and TEM20 without unlock the cavity as before.
| ||AOM Offset [V] ||SC Peak Amplitude [V] ||BAB Peak Amplitude [V] |
|TEM00 ||0.0008 ||1 ||0.16 |
|TEM10/TEM01 ||-3.623 ||0.18 ||0.005 |
|TEM20/02 ||3.475 ||0.125 ||1e-3 |
At the end of this analysis we fond that the SC was no more perfectly co-resonant with the BAB. We moved again it PLL offset to 100.09417962 MHz.
With the double pass the Green/IR coresonant is less stable, thus we need to refine oftenly the AOM offset.
Green Lock trigger:
We worked on the lock trigger of the FC on the green beam in order to lock in the green peak corresponding to the IR peak. To do this in the LFC_Acl process we set the thresholds of the SC_PD_DC_CND realy and we added it into the lock filter reset condition.
We tested it but it is not working well. In particular if the AOM offset is not perfectly tuned the cavity does not lock anymore on the green beam. Moreover if we leave the cavity unlocked for more than 5 minutes the misaligment degradates the IR/Green coresonance and the trigger does not work anymore.
We disabled it by focing the relay to stay alwais at 1.
Green EOM modulation frequency back:
- From 22:17:30 UTC stretch of data with the GR EOM frequency set to 5.52947075MHz
- From 22:34:30 UTC stretch of data wih the GR EOM frequency set to 5.52942697MHz (wrong demod phase 0.1 rad)
- From 22:47:30 UTC stretch of data wih the GR EOM frequency set to 5.52942697MHz (right demod phase 0.22 rad)
During this test the cavity was locked on green, the IR was resonant and the SC was on.
Week end data acquisition:
Before leaving I saved all the AA mirror set points in the various configuration files.
We leave the system in the following way
- Green EOM freuquency restored at the initial value
- AOM scan between -0.04 and 0.08V at 0.1Hz
- BAB and SC distance of about 1 kHz SC PLL offset: 100.09517963 MHz
The AOM ramp was switched on at 23:01 UTC
Data will be analyzed